# Identifying and Mitigating API Misuse in Large Language Models

### v1

@@ 3.2 Evaluation Elements
To invoke an API correctly, LLMs are expected to predict both the correct API method name and pass valid arguments for the API parameters. We design the following two scenarios to evaluate the correctness of the generated API elements:
API Method Inference
In this evaluation setup, we mask out the API method name in a code snippet and ask the models to infer the appropriate method, similar to type inference [31, 32]. This evaluation assesses whether the models can comprehend the intent of the code context and recommend a suitable method name based on the imported packages and the parameters’ context.
@@
Figure 2: Example for API Method Inference: The model should predict ‘get‘ as the appropriate method name.
@@
API Parameter Generation
In this case, rather than focusing on predicting the method name, we evaluate the models’ ability to generate valid and contextually appropriate parameter values for a given API method. Here, the model is provided with a partially completed code snippet where the API method name is already specified, but all the context (except for the opening parenthesis) is removed. The models are expected to complete the arguments correctly by recalling the parameter names and types associated with the API method and identifying potential variables as valid arguments.
@@
Figure 3: Example for API Parameter Generation: The model should predict ‘url’ as the appropriate parameter.
@@
3.3 LLM Models
LLMs typically differ in their architecture, size, and pretraining tasks. In practice, decoder-only models are preferred in the IDE environment, where developers utilize models to complete code and generate comments [33]. They may also commonly use the LLM to summarise code to aid understanding and to generate test cases for target code. Current IDEs like VSCode and Cursor [^3] https://www.cursor.com/ provide two types of decoder-only models for users: Open-source LLMs and Closed-source LLMs. Open-source LLMs are typically deployed in the user’s environment to optimize the inference speed (e.g., low latency) and flexibility (e.g., domain-specific fine-tuning) [34], while closed-source LLMs are more powerful but can only be inferred via company-hosted remote APIs. To study and compare representative examples of these different LLM models, we choose the widely used open-source model, StarCoder, and Copilot, a closed-source model provided by GitHub.
@@ 3.3 LLM Models
Copilot is a closed-source code suggestion model based on OpenAI’s GPT-4, initially launched in 2022. While the model is further fine-tuned for coding tasks, it has a much better understanding of natural language context than open-source LLMs like StarCoder. More than 46% of developers’ code files, on average, are reportedly generated by GitHub Copilot. Similar to StarCoder, Copilot offers two generation paradigms: code completion and code infilling. Besides, Copilot utilizes a client-side model to reduce the frequency of unwanted suggestions when they might prove disruptive to a developer’s workflow, helping it better respond to each developer using it. In this work, we use the Copilot version, which started in 2024, and manually paste code snippets into VSCode for model inference.
@@ 3.4 Hyperparameters
To align real-world software development in the IDE with LLMs, we use the default hyperparameters specified in each VSCode extension. Regarding StarCoder, we follow the configurations of Hugging Face’s llm-ls [^4] https://github.com/huggingface/llm-ls, a Language Server Protocol server leveraging LLMs for code completion. Specifically, we use the temperature of 0.2, the maximum new token number of 150, and the top_p value of 0.95. For GitHub’s Copilot model, we use the official v1.252.0 extension in VSCode 1.95.3 and prompt the model with the code snippet in a single file. Specifically, we directly query the models for autocompletion without any customized instructions.
@@ 3.5 Dataset
We describe how we construct the dataset for our qualitative study. We leverage The Stack [35] data, the largest and most up-to-date permissively licensed GitHub corpus containing source code files in 358 programming languages from GitHub repositories. Due to the high data quality resulting from the comprehensive filtering mechanisms, The Stack dataset has been partially utilized by a series of LLMs for pretraining. Specifically, we use v1.2, the latest version that opted out of the requested and malicious data, and exclude the subset used to train StarCoder to avoid data contamination issues [36].
To conduct a thorough investigation, we selected two popular programming languages, Python and Java. As each code file can contain multiple library APIs, we localize all the positions of APIs (methods and parameters) via static analysis. It is important to note that there could be alternative signatures for the same APIs, particularly where methods share the same name but differ in their parameters (e.g., overloading or different argument types). Additionally, we will consider APIs that have the same function or call methods but are located in different paths or namespaces due to different versions. To account for this, we ensure that all distinct signatures are included in our dataset. For example, an API could appear in various forms depending on the number and types of parameters used. To ensure comprehensive coverage of different programming scenarios, we include only one sample for each fully qualified-named API signature in our dataset, which helps us capture the diversity of API usage across various contexts. Furthermore, there are many duplicated APIs invoked in various code snippets. To ensure that our examination can cover diverse programming scenarios in the wild, we choose only one sample for each unique, fully-qualified, named API signature for each file. In total, we include 3,000 samples for both Python and Java, respectively, based on the most frequently used APIs. For each sample of API usage, we treat the code before the API method as a prefix and the parenthesized parameters as a suffix for Parameter Prediction and the code before the API parameter as a prefix for Parameter Generation, which results in 24,000 generated samples.
@@ 3.6 Evaluation Metrics
As we randomly sample our dataset from a raw GitHub code corpus without environment configurations (e.g., dependencies) and oracles, it is hard to assess the code quality via compilation and test cases. Following the Prior works [37, 38, 39], we use BLEU [40], CodeBERTScore [41], and Exact Match (EM) as the evaluation metrics to determine the (un)likelihood that the models can invoke the same APIs as human developers.
BLEU
BLEU calculates the percentage of 4-gram overlap between the reference fixed code and the candidate code generated by the models. For simplicity, we refer to it as BLEU in our work. The candidate fixed code represents the code generated by the models, while the reference fixed code represents the ground-truth code written by developers. BLEU is defined as follows:
@@
where $p_{n}$ is the precision of n-grams, $w_{n}$ is the weight for n-grams (usually equal), and BP is the brevity penalty to account for shorter candidate sequences.
CodeBERTScore
CodeBERTScore, inspired by BERTScore [42], measures semantic consistency between generated and reference code by computing the cosine similarity of their token representations. Specifically, CodeBERTScore computes the F scores by combining the precision and recall. Precision ($\text{CodeBERTScore}_{P}$) is calculated as the average maximum similarity between each token in the generated code and tokens in the reference code, while recall ($\text{CodeBERTScore}_{R}$) is computed as the average maximum similarity between each token in the reference code and tokens in the generated code. These are combined using a weighted harmonic mean to produce the final score. In this work, we adopt $\text{CodeBERTScore}_{F_{3}}$, which prioritizes recall over precision, as it better aligns with functional correctness in code generation tasks. The formula is defined as:
@@
Exact Match (EM)
EM is defined as the count of generated code instances $\hat{Y}$ whose abstract syntax trees (ASTs) exactly match the ASTs of the corresponding reference/ground-truth code $Y^{*}$. It is a strict metric that evaluates syntactic equivalence and may exclude reasonable outputs that differ in syntax but are semantically correct. EM serves as a lower bound, as different programs can achieve the same functionality but be written differently. In this context, $\hat{Y}$ represents the set of generated code instances, and $Y^{*}$ represents the set of reference or ground-truth code instances existing in The Stack dataset, which well captures how humans invoke those APIs. Each $\hat{y}$ is a single generated code instance, and each $y^{*}$ is a corresponding reference code instance. The EM metric counts the number of AST-based exact matches in the test set instead of the exact values passed into each API. The formula for EM is defined as:
@@ 3.7 EM Failure Qualitative Example Analysis
Two authors, each with over five years of programming experience in Python and Java, reviewed and categorized the API misuse samples. We apply open coding to label the errors descriptively without predefined categories. Through constant comparison, they iteratively grouped similar misuses into higher-level conceptual categories.
First, the two authors independently familiarized themselves with the dataset by reviewing a subset of API misuse cases. Each annotator labeled 50 randomly selected samples, took detailed notes on their observations, During this process, they were allowed to search online for API information when necessary to validate the correctness of the method and parameter usage. Following this initial review, the annotators conducted axial coding, systematically grouping similar misuses into broader conceptual categories based on shared characteristics. They then met to compare their annotations, resolve discrepancies, and iteratively refine their categorization framework. Through multiple review cycles, they refined their labels, clarified ambiguous definitions, and introduced new categories where necessary. In the end, they progressively consolidated similar misuse patterns into four distinct categories: Intent Misuse, Hallucination, Missing Parameter, Redundant API Call. With the finalized codebook, the two authors independently labeled the remaining API misuse cases. To assess inter-annotator agreement, we computed Cohen’s Kappa score [43], achieving a final value of 0.97, indicating almost perfect agreement. The annotation process spanned 200 hours, encompassing multiple iterations of refinement to ensure accuracy and consistency.
@@
- absolute_distances = np.abs(to_points - from_point)
@@
Figure 4: A diff-formatted example of intent misuse: Incorrect use of np.abs instead of vx.magnitude in Python.
@@ 3.8 API Misuse Classification
Type 1: Intent Misuse
Intent misuse occurs when a model selects an API or method that is syntactically correct but does not align with the intended functionality or task. This leads to incorrect behavior due to the model’s misunderstanding of the method’s purpose. In Figure 4, the method farthest() is intended to find the farthest point from a reference point. To calculate this, the absolute distances between points must be computed. The model incorrectly uses np.abs, a method from the numpy library, which computes the element-wise absolute values of an array. Although syntactically correct, np.abs does not calculate the magnitude of the vectors, which is the intended purpose in this case. The correct API is vx.magnitude, which is designed to compute the magnitude of vectors and is thus more suitable for the task. Therefore, the misuse here stems from the model’s failure to select the appropriate method that aligns with the intended operation, resulting in a mismatch between the code’s functionality and its purpose.
@@
- parameterPolynomial.setCubic(trajectoryTime.getDoubleValue(), 0.0, Double.POSITIVE_INFINITY);
@@
Figure 5: Diff-formatted Examples of hallucination misuse: navigation_main_menu_titles instead of navigation_main_titles and non-existent setCubic instead of checkIntervalContains in Java.
Type 2: Hallucination Misuse
Hallucination occurs when a model introduces an entirely incorrect API method or parameter that does not exist or is not required for the task. The model may create or "hallucinate" an API call that seems plausible but invalid in the context, leading to erroneous behavior or a compilation/runtime error. Concurrently, several studies have studied hallucination issues of LLM-based code generation [44, 45, 46]. However, they consider any bugs made by LLMs to be a type of hallucination. In this work, we define “Hallucination” as the case when LLMs generate non-existent API parameters or methods in the wild. In Figure 5, the model incorrectly uses the method res.getStringArray(R.array.
navigation_main_menu_titles)
as a parameter for the Collections.addAll() method. The correct call should be res.getStringArray(R.array.
navigation_main_titles)
, where the parameter name was likely fabricated by the model, leading to an API misuse. This is a case of parameter hallucination, where the model introduces a nonexistent or incorrect parameter in the API call. In Figure 5, the model wrongly selects parameterPolynomial.setCubic(trajectoryTime.
getDoubleValue(), 0.0, Double.POSITIVE_INFINITY)
when the correct method should have been
MathTools.checkIntervalContains() for checking a time interval. Here, the model replaces the correct API method with a non-existent or irrelevant one, resulting in method hallucination. The model mistakenly introduces an incorrect method that does not align with the context or expected API behavior, leading to a significant deviation from the intended functionality.
@@
Type 3: Missing Item Misuse
@@
Figure 7: A diff-formatted example of redundancy misuse: Unnecessary chaining of initialize() after loadNow() in Java. The correct implementation should avoid redundant method calls where a single method is sufficient.
Type 4: Redundancy

### v2

@@ 3.2 Evaluation Elements
To invoke an API correctly, LLMs are expected to predict both the correct API method name and pass valid arguments for the API parameters, where the objectives are similar to the human-written code [schlichtig2022fum]. We design the following two scenarios to evaluate the correctness of the generated API elements:
API Method Infilling
In this evaluation setup, we mask out the API method name in a code snippet and ask the models to infer the appropriate method, similar to type inference [cassano2023type, peng2023generative]. This evaluation assesses whether the models can comprehend the intent of the code context and recommend a suitable method name based on the imported packages and the parameters’ context.
@@
Figure 2: Example for API Method Infilling: The model should predict ‘get‘ as the appropriate method name.
@@
API Parameter Completion
In this case, rather than focusing on predicting the method name, we evaluate the models’ ability to generate valid and contextually appropriate parameter values for a given API method. Here, the model is provided with a code snippet where the API method name is already specified, and it is tasked with generating the appropriate argument(s) within the method call. All surrounding context after the method name is masked, requiring the model to infer correct usage based on its knowledge of the API. The models are expected to complete the arguments correctly by recalling the parameter names and types associated with the API method and identifying potential variables as valid arguments.
@@
Figure 3: Example for API Parameter Completion: The model should predict ‘url’ as the appropriate parameter.
@@
For evaluation, Exact Match (EM) is computed at the level of the masked code elements (API method name or parameters), rather than the entire snippet, to directly capture API correctness. In addition to BLEU and CodeBERTScore, we also discuss precision and recall at the API-element level to better reflect the accuracy of API invocations.
3.3 Models
LLMs differ in architecture, size, and pretraining tasks. In practice, decoder-only models are preferred in IDE environments, where developers rely on them for code completion and comment generation [svyatkovskiy2020intellicode]. They are also used to summarize code for comprehension or generate test cases for target snippets. Current IDEs such as VSCode and Cursor [^4] https://www.cursor.com/ primarily deploy decoder-only base models that directly predict the next tokens without instruction tuning. This design supports fast inference and seamless integration into interactive editing workflows [izadi2024language, takerngsaksiri2024students, semenkin2025full]. Open-source LLMs are typically executed locally to maximize inference speed and customization flexibility (e.g., domain-specific fine-tuning) [fu2024serverlessllm], while closed-source LLMs offer higher performance but can only be accessed via hosted APIs. To capture representative behaviors of LLMs commonly used in IDEs, we began our experiments in early 2024 and selected three widely used models for code generation [Wiggers2023StarCoderFreeModel, hui2024qwen2, IBM2023watsonxMainframe, Dohmke2023CopilotTransformsPlatform]: StarCoder-7B, an open-source decoder-only model from Hugging Face; Qwen2.5-Coder-7B, an open-source model from Alibaba’s Qwen2.5 family; and GitHub Copilot, a closed-source model powered by OpenAI’s GPT-4. These models collectively represent both open and closed ecosystems, balancing between transparency, accessibility, and real-world deployment in IDE workflows. Although newer versions such as StarCoder2 have since been released, our selected models remain representative of IDE-integrated LLMs at the time of study.
@@ 3.3 Models
Qwen2.5-Coder-7B is a recent open-source code generation model from Alibaba’s Qwen2.5 series. Like StarCoder, it adopts a decoder-only transformer architecture optimized for next-token and fill-in-the-middle prediction tasks. It is pretrained on a large corpus of multilingual source code and natural language, supporting over 90 programming languages. Compared with StarCoder, Qwen2.5-Coder incorporates more recent repositories and improved tokenizer coverage for Python and Java, making it well-suited for IDE-style code completion without requiring instruction tuning.
Copilot is a closed-source small code suggestion model distilled from OpenAI’s GPT-4, initially launched in 2022. While the model is further fine-tuned for coding tasks, it has a much better understanding of natural language context than open-source LLMs like StarCoder. More than 46% of developers’ code files, on average, are reportedly generated by GitHub Copilot. Similar to StarCoder, Copilot offers two generation paradigms: code completion and code infilling. Besides, Copilot utilizes a client-side model to reduce the frequency of unwanted suggestions when they might prove disruptive to a developer’s workflow, helping it better respond to each developer using it. In this work, we use the Copilot version, which started in 2024, and manually paste code snippets into VSCode for model inference.
@@ 3.4 Hyperparameters
To align real-world software development in the IDE with LLMs, we use the default hyperparameters specified in each VSCode extension. Regarding StarCoder, we follow the configurations of Hugging Face’s llm-ls [^5] https://github.com/huggingface/llm-ls, a Language Server Protocol server leveraging LLMs for code completion. Specifically, we use the temperature of 0.2, the maximum new token number of 150, and the top_p value of 0.95. For Qwen2.5-Coder-7B, we use the recommended hyperparameters, with the temperature of 0.3 and the top_p value of 0.95. For GitHub’s Copilot model, we use the official v1.252.0 extension in VSCode 1.95.3 and prompt the model with the code snippet in a single file. Specifically, we directly query the models for autocompletion without any customized instructions.
@@ 3.5 Dataset
We describe how we construct the dataset for our qualitative study. We leverage The Stack [kocetkovstack] data, the largest and most up-to-date permissively licensed GitHub corpus containing source code files in 358 programming languages from GitHub repositories. Due to the high data quality resulting from the comprehensive filtering mechanisms, The Stack dataset has been partially utilized by a series of LLMs for pretraining. Specifically, we use v1.2, the latest version that opted out of the requested and malicious data, and exclude the subset used to train StarCoder to avoid data contamination issues [sainz2023nlp].
To conduct a thorough investigation, we selected two popular programming languages, Python and Java. As each code file can contain multiple library APIs, we localize all the positions of APIs (methods and parameters) via static analysis. It is important to note that there could be alternative signatures for the same APIs, particularly where methods share the same name but differ in their parameters (e.g., overloading or different argument types). Additionally, we consider APIs that provide the same functionality but reside in different paths or namespaces due to version differences. To account for this, we ensure that all distinct signatures are included in our dataset. For example, an API could appear in various forms depending on the number and types of parameters used. To ensure comprehensive coverage of different programming scenarios, we include only one sample for each fully qualified-named API signature in our dataset, which helps us capture the diversity of API usage across various contexts. Furthermore, there are many duplicated APIs invoked in various code snippets. To ensure our examination covers diverse programming scenarios in the wild, we restrict the dataset to at most one instance of each unique, fully-qualified API signature per file. This prevents highly frequent APIs (e.g., numpy.array, requests.get) from dominating the dataset while still allowing us to capture multiple usages of the same API across different projects and contexts. In total, we randomly sample 3,000 unique APIs for both Python and Java, respectively, based on their distinct fully qualified names. The APIs are collected from 631 Python and 486 Java repositories, covering 4,641 Python packages and 6,258 Java packages identified through import statements. This broad collection ensures coverage of diverse libraries and API usage contexts, reflecting a wide range of real-world development practices. For each API usage sample, we treat the code before the API method as a prefix and the parenthesized parameters as a suffix for API Method Infilling, and the code before the API parameter as a prefix for API Parameter Completion, resulting in 36,000 generated samples.
Although The Stack applies rigorous filtering, we acknowledge that human-written code may still contain errors; therefore, we treat it as an approximate ground truth, assuming that the majority of examples represent correct API usage while recognizing that occasional misuses or stylistic inconsistencies may introduce noise. Since our dataset emphasizes frequently used APIs, the findings may not fully generalize to rarely used or newly introduced APIs, which we discuss further in threats to validity in subsection 6.3.
@@ 3.6 Evaluation Metrics
As we randomly sample our dataset from a raw GitHub code corpus without environment configurations (e.g., dependencies) and oracles, it is hard to assess the code quality via compilation and test cases. Following prior works [hu2024active, liu2023codegen4libs, ma2024compositional], we use BLEU [papineni2002bleu], CodeBERTScore [zhou2023codebertscore], and Exact Match (EM) as evaluation metrics to determine the (un)likelihood that the models can invoke the same APIs as human developers. Since automatic metrics have known limitations, we further complement them with element-level precision and recall, as well as manual analysis of misuse categories.
BLEU BLEU calculates the percentage of 4-gram overlap between the reference fixed code and the candidate code generated by the models. For simplicity, we refer to it as BLEU in our work. The candidate fixed code represents the code generated by the models, while the reference fixed code represents the ground-truth code written by developers. BLEU is defined as follows:
@@
where $p_{n}$ is the precision of n-grams, $w_{n}$ is the weight for n-grams (usually equal), and BP is the brevity penalty to account for shorter candidate sequences. BLEU captures overall fluency but may be diluted by tokens outside the masked API elements.
CodeBERTScore CodeBERTScore, inspired by BERTScore [zhangbertscore], measures semantic consistency between generated and reference code by computing the cosine similarity of their token representations. Specifically, CodeBERTScore computes the F scores by combining the precision and recall. Precision ($\text{CodeBERTScore}_{P}$) is calculated as the average maximum similarity between each token in the generated code and tokens in the reference code, while recall ($\text{CodeBERTScore}_{R}$) is computed as the average maximum similarity between each token in the reference code and tokens in the generated code. These are combined using a weighted harmonic mean to produce the final score. In this work, we adopt $\text{CodeBERTScore}_{F_{3}}$, which prioritizes recall over precision, as it better aligns with functional correctness in code generation tasks. The formula is defined as:
@@
While CodeBERTScore is more flexible than EM in capturing semantic similarity, it has limitations. In particular, it may underestimate correctness when two different but functionally equivalent APIs are used, since token-level embeddings may not capture equivalence unless such relationships were observed during pretraining. Thus, CodeBERTScore should be interpreted as an approximation of semantic similarity.
Exact Match (EM) EM is defined as the count of generated code instances $\hat{Y}$ whose abstract syntax trees (ASTs) exactly match the ASTs of the corresponding reference/ground-truth code $Y^{*}$. It is a strict metric that evaluates syntactic equivalence and may exclude reasonable outputs that differ in syntax but are semantically correct. In our study, EM is computed at the level of the masked API element (method name or parameter), rather than the entire snippet, to directly assess correctness of API usage. EM serves as a lower bound, as different programs can achieve the same functionality but be written differently. The formula is defined as:
@@
A higher EM score consistently indicates stronger performance. However, since human-written code is used as ground truth, EM may penalize models that generate different but still valid API usages. We therefore report EM alongside BLEU, CodeBERTScore, and precision/recall to provide a more complete evaluation of API invocation correctness.
@@ 3.7 EM Failure Qualitative Example Analysis
Two authors, each with over five years of programming experience in Python and Java, reviewed and categorized the API misuse samples. We adopt a closed coding strategy by mapping LLM-generated API misuses to Wei et al.’s taxonomy [wei2024demystifying], which classifies misuses along two dimensions: violation types (Missing, Redundant, Replacement, Outdated) and API elements (Method, Parameter, Condition). However, because our evaluation is based on isolated code snippets, without access to full projects, temporal version histories, or documentation, we focus only on the violation types that are reliably annotatable in this setting: Missing and Redundant.
We exclude Replacement and Outdated categories, as accurately identifying whether a method has been superseded (Outdated) or semantically replaced (Replacement) often requires knowledge of API evolution, project-specific context, or historical intent, which are unavailable in our static snippet-based setting. For example, detecting an outdated method would require knowing the release timeline or deprecation status of the corresponding library, while identifying a replacement misuse would require understanding the intended functionality and its modern equivalent. Since these contextual signals are absent from isolated code samples, including them could lead to annotation ambiguity and unreliable labeling. Consequently, we focus on misuses that can be confidently recognized from self-contained code evidence while maintaining conceptual consistency with Wei et al.’s taxonomy [wei2024demystifying].
To analyze LLM-specific characteristics, we annotate two recurring surface patterns, intent misuse and hallucination, as tags within the broader Missing/Redundant framework. For example, hallucinated methods typically reflect a non-existent method generated in place of a missing or misused one, and can be understood as a form of Missing Method. Similarly, intent misuses, while superficially aligned with Replacement–Method, are treated as semantic misfires under Redundant or Incorrect Method, when replacement judgment is unreliable. This constrained coding preserves alignment with prior taxonomies while recognizing the limits of static code-only analysis in LLM-generated contexts.
First, the two authors independently familiarized themselves with the dataset by reviewing a subset of API misuse cases. Each annotator labelled 50 randomly selected samples, took detailed notes on their observations, During this process, they were allowed to search online for API information when necessary to validate the correctness of the method and parameter usage. Following this initial review, the annotators conducted axial coding, systematically grouping similar misuses into broader conceptual categories based on shared characteristics. They then met to compare their annotations, resolve discrepancies, and iteratively refine their categorization framework. Through multiple review cycles, they refined their labels, clarified ambiguous definitions, and introduced new categories where necessary. In the end, they progressively consolidated similar misuse patterns into four distinct categories: Intent Misuse, Hallucination, Missing Parameter, Redundant API Call. With the finalized codebook, the two authors independently labelled the remaining API misuse cases. To assess inter-annotator agreement, we computed Cohen’s Kappa score [mchugh2012interrater], achieving a final value of 0.97, indicating almost perfect agreement. The annotation process spanned 300 hours, encompassing multiple iterations of refinement to ensure accuracy and consistency.
@@
- absolute_distances = vs.abs(to_points - from_point)
@@
Figure 4: A diff-formatted example of intent misuse: Incorrect use of abs instead of magnitude in Python.
@@ 3.8 API Misuse Classification
LLM-Specific – Intent Misuse
Intent misuse occurs when a model selects an API or method that is syntactically correct but does not align with the intended functionality or task. This results in incorrect behavior due to the model’s misunderstanding of the method’s purpose. In Figure 4, the method farthest() is intended to compute the farthest point from a reference point, which requires calculating vector magnitudes. The model incorrectly uses np.abs, a numpy function that returns element-wise absolute values, rather than vx.magnitude, which directly computes vector magnitudes. Although syntactically valid, this misuse reflects a semantic mismatch between the selected API and the intended operation. While such cases may be superficially mapped to the Replacement–Method category in Wei et al.’s taxonomy [wei2024demystifying], prior studies have not systematically examined intent misuses in the context of LLM-generated code. We adapt and extend the taxonomy to explicitly capture this recurring pattern, which emerges frequently in LLM-assisted code completion. Importantly, we distinguish intent misuses from cases where multiple APIs could plausibly achieve the same functionality (e.g., different logging or plotting methods). In such cases, annotators did not label the output as misuse, since the generated API remained consistent with the overall intent of the code.
@@
- MathTools.setCubic(trajectoryTime.getDoubleValue(), 0.0, Double.POSITIVE_INFINITY);
@@
Figure 5: Diff-formatted Examples of hallucination misuse: non-existent setCubic instead of checkIntervalContains in Java.
LLM-Specific – Hallucination Misuse
Hallucination occurs when a model introduces an entirely incorrect API method or parameter that does not exist or is not required for the task. The model may fabricate an API call that seems plausible but is invalid in context, leading to erroneous behavior or compilation/runtime errors. Although prior work has discussed hallucination broadly in LLMs [liu2024exploring, zhang2024llm, tian2024codehalu], this work presents the first systematic categorization of hallucination as an API misuse type, where models generate non-existent API methods or parameters. We adapt existing taxonomies by explicitly introducing this category to capture a recurring and distinctive error pattern in LLM-generated code. As shown in Figure 5, the model incorrectly predicts res.getStringArray(R.array.navigation
_main_menu_titles)
instead of the correct res.getStringArray(R.array.navigation
_main_titles)
. The fabricated parameter name represents a case of parameter hallucination. In the same figure, the model generates a non-existent method call parameterPolynomial.setCubic(...) in place of the valid MathTools.checkIntervalContains(), illustrating method hallucination. Both cases highlight how LLMs may generate plausible but invalid APIs, a failure mode that is not emphasized in prior human-centric misuse studies but emerges prominently in code completion tasks.
@@
Generic – Missing Item Misuse [wei2024demystifying]
@@
Figure 7: A diff-formatted example of redundancy misuse: Unnecessary chaining of initialize() after loadNow() in Java. According to the official API documentation, loadNow() already performs the necessary initialization, so an additional call to initialize() has no effect and is therefore redundant.
Generic – Redundancy [wei2024demystifying]